# Uplifting Lower-Income Data: Strategies for Socioeconomic Perspective Shifts in Vision-Language Models

### v2

@@ Abstract
Unequal representation across cultures and socioeconomic groups in AI is a significant and challenging problem, often leading to uneven model performance. As a step toward addressing this issue, we formulate translated non-English, geographic, and socioeconomic integrated prompts and evaluate their impact on VL model performance for data from different countries and income groups. Our findings show that geographic and socioeconomic integrated prompts improve VL performance on lower-income data and favor the retrieval of topic appearances commonly found in data from low-income households. From our analyses, we identify and highlight contexts where these strategies yield the most improvements. Our model analysis code is publicly available at Analysis for Uplifting lower-income data.
@@
Socioeconomic Perspective Shifts in Vision-Language Models
@@ 1 Introduction
The overrepresentation of Western data and lack of diversity are common issues in many popular datasets Shankar et al. (2017). Even though the size and quality of pre-training datasets greatly impact the performance of today’s AI models, there is insufficient research attention given to this area. Furthermore, there is even less attention given to data curation methods that can improve representation in AI datasets Sambasivan et al. (2021).
Today, vision-language (VL) models are leveraged to filter uncurated data repositories into training datasets by assessing the association strength between images and text Fang et al. (2023). For example, Open AI’s ViT-B/32 Radford et al. (2021) was used to filter web-scraped images to create the LAION-5B dataset Schuhmann et al. (2022). However, many VL models like CLIP have been shown to perform unequally for data from different cultures and socioeconomic classes Nwatu et al. (2023).
Since datasets filtered using this technique reflect the VL model used for filtering Fang et al. (2023), this practice can exacerbate the lack of diversity in AI models and datasets. This is evident in Ignat et al. (2024), which shows that data from the LAION-5B dataset is most similar to data from Western countries, like the United States and Canada, while it is dissimilar to data from many non-Western countries.
To mitigate the lack of representation in AI, we address the issue of performance inequality in VL models Radford et al. (2021); Visheratin (2023) through prompting that leverages the cultural knowledge embedded in language Ventura et al. (2023); Buettner et al. (2024); Nguyen et al. (2024). Our goal is to improve the performance of VL models on the diverse representations of topic labels especially found in data from households with non-Western and lower socioeconomic status. Specifically, we pose several research questions to evaluate the effects of non-English languages and geographic and socioeconomic attribute-integrated prompts on retrieving diverse images. We focus on identifying which strategies improve performance on lower-income data, while also noting their effects on higher-income data.
Our contributions are summarized as follows. First, we show that prompts translated to the native language (referred to as the non-English major language) of a country currently do not yield improved Recall performance on images from that country compared to English prompts. Second, by conducting an in-depth analysis of vision-language models’ understanding of these attributes and their effects on Recall across data from different countries, we establish that geographic attribute and socioeconomic attribute integrated prompts improve VL performance on lower-income data, and identify contexts where these prompts work best. Third, we share insights from our analysis demonstrating how these attributes drive a perspective shift that benefits the retrieval of lower-income data.
@@ 2 Related Work
Class imbalances in training data contribute greatly to bias propagation in AI models Ferrara (2024); Shankar et al. (2017); He and Garcia (2009); Pouget et al. (2024). This bias manifests in the disparate impact of these models on users in applications such as facial recognition Buolamwini and Gebru (2018), healthcare Obermeyer et al. (2019), and hiring Raghavan et al. (2020). Since creating balanced diverse datasets is difficult and expensive Ignat et al. (2024); Ramaswamy et al. (2023), the research community has explored alternative methods of dealing with bias. Some of these methods involve pre-processing and in-processing techniques such as data augmentation, feature importance tuning, regularization, and adversarial training Yan et al. (2020); Zafar et al. (2017); Ignat et al. (2024); Maudslay et al. (2019); Sharma et al. (2020); Navarro et al. (2024); Zhang et al. (2018). However, our work is most similar to post-processing techniques such as Ferrara (2023); Hardt et al. (2016); Kamiran et al. (2012); Pleiss et al. (2017) that adjust model outcomes using a set of criteria to fit diversity standards across race, gender, and culture for the benefit of disadvantaged groups. Pouget et al. (2024); Nwatu et al. (2023) show that vision-language models consistently perform badly on data from lower socioeconomic status. Our analysis seeks to identify and analyze the effects of non-invasive post-processing techniques that mitigate this issue.
@@
Language is often referred to as a vehicle for propagating cultural knowledge and norms Callies (2024); Sharifian (2014); Karsdorp and Fonteyn (2019); Norton (1997). This is exemplified in AI research, where models often pick up on biases contained in the language of their training data Stanczak and Augenstein (2021); Rogers et al. (2021), and model outputs can be controlled by specifying a cultural shift in perspective Ventura et al. (2023) to improve diversity.
Furthermore, Arora et al. (2023); Cao et al. (2023); AlKhamissi et al. (2024); Liu et al. (2021) have shown that while LLMs and vision-language models contain a decent amount of cultural information regarding cultures present in English data (predominantly Western), for cultures contained in non-English language data, not as much cultural information can be extracted. Major reasons for this disparity include differences in the quantity and quality of training data available for these languages compared to English, information loss due to language translation, and model design decisions Arora et al. (2023); Hershcovich et al. (2022); Nasif et al. (1991).
Similar to De Vries et al. (2019) and Nguyen et al. (2024) that demonstrate how language can improve data diversity by retrieving images using translated captions, our work seeks to identify how current multilingual VL models and non-English languages can improve representation in vision-language models and datasets across regions and income groups.
@@
Prompting techniques for large language models have been extensively studied in recent years. Both hard prompting Petroni et al. (2019); Zhou et al. (2023) and soft prompting Huang et al. (2023); Goswami et al. (2023) are useful in adapting models for downstream tasks, instruction tuning, and value alignment. Prompts have also been used in vision-language models for similar purposes Lu et al. (2022); Yao et al. (2024); Zhou et al. (2022). Most similar to our work, Buettner et al. (2024) incorporates geographic and physical attributes of objects into prompts to improve retrieval of diverse images. However, we extend the investigation to non-English language prompts and socioeconomic attributes and then perform further analysis to highlight insights into how VL models encode representations of different topics across not only regions but also socioeconomic status.
@@ 3 Methodology
We apply three types of text prompting techniques to a geographically and socio-economically diverse dataset and analyze how these changes affect the performance of a multilingual VL model on data across different socio-economic groups, especially focusing on lower-income data.
@@ 3.1 Dollar Street Dataset
Income Classes.
We categorize the images and countries on Dollar Street based on their income information as described below.
@@
For fair comparison across bins, we group the images using the quartile binning method which splits the data into an approximately equal number of images per bin as shown in Rojas et al. (2022). We group the images into four income classes (“poor”, “low-mid”, “up-mid”, and “rich” ) using quartiles as shown in Table 1. We further categorize the lowest two image income classes as lower-income images and the highest two income groups as higher-income images.
@@
Country Income Classes.
We group all the 63 countries from Dollar Street into country income classes based on their World Bank income classification.111https://datahelpdesk.worldbank.org/ All the countries and their income classes are shown in Section A.2. We further categorize the lowest two country income classes as lower-income countries and the highest two income groups as higher-income countries.
@@
There are unique topics that reflect common household objects and human actions (e.g., “toilet paper”, “get water”), some of which are subjective (e.g., “next big thing I plan to buy”, “favorite sports clubs”, “most loved item”). Following De Vries et al. (2019) and Nwatu et al. (2023), we remove nineteen subjective topics from the dataset.
@@ 3.2 Prompt Design
Specifically, we pair each country with their non-English major language (e.g., Spanish for Brazil, French for Cameroon) following the country and language information provided by official sources.222https://www.cia.gov/the-world-factbook/field/languages/, https://www.ncsc.org/__data/assets/pdf_file/0024/17862/languagesbycountries.pdf, https://www.dss.gov.au/sites/default/files/files/foi_disclosure_log/12-12-13/language-list.pdf
We identify 59/63 countries in Dollar Street where one or more non-English major languages are spoken. We also select languages covered by state-of-the-art machine translation and multilingual vision-language models. There are 40 such non-English major languages, and they are listed in Section A.2.
Finally, we translate the default English prompts to these 40 languages using the NLLB-200-distilled-600M Costa-jussà et al. (2022), a state-of-the-art neural machine translation model. If an image prompt is translated into the non-English major language of the image’s country of origin, it is referred to as a native translated prompt.
@@
3.3 State-of-the-art Vision-Language Model
For our evaluation, we chose NLLB-CLIP-SigLIP Visheratin (2023), a state-of-the-art multilingual vision-language model, due to its wide reach across many low-resource languages and superior performance among other models.333https://huggingface.co/visheratin/nllb-clip-large-siglip The model consists of an image encoder from the SigLIP model Beyer et al. (2022); Zhai et al. (2023) and a text encoder from the NLLB model Costa-jussà et al. (2022). The model supports the 201 languages of the Flores-200 Costa-jussà et al. (2022) and has recorded groundbreaking results on the Crossmodal-3600 dataset Thapliyal et al. (2022), especially on low-resource languages.
@@ 4 Research Questions
We perform several analyses to answer three research questions that uncover and mitigate limitations in the performance of VL models across different countries and socioeconomic groups.
@@
RQ1. Do translated prompts improve performance for lower-income images?
@@
We calculate the cosine similarities between the NLLB-CLIP-SigLIP image embeddings and the text embeddings of the translated prompts for each image and topic pair and the 40 non-English languages. This process yields 40 image-topic alignment scores for each image across all non-English languages.
@@
As a baseline, we compute the alignment scores between the images and the default English prompts. We use the alignment scores to compute the Recall scores for each topic. Specifically, we select the top N images with the highest alignment scores for the given topic, where N represents the number of ground truth images corresponding to that topic. We group and analyze the Recall scores across different countries and image income classes and present our findings below.
Native translated prompts perform consistently worse than English prompts on lower-income images from their respective countries.
@@
We focus our analysis on images from the two lowest image income groups, i.e., poor and low-middle. We filter out countries with no data for these income groups (e.g., Russia, Turkey) and obtain 39/59 countries and 28/40 non-English languages. We pair up each of the remaining countries with the results from their respective non-English native languages and compare the Recall scores obtained using the native translated prompts with those obtained using the default English prompts.
We find that native translated prompts consistently perform worse than default English prompts. This is shown in Figure 2, where we compute the average Recall across all countries for English and the native translated prompts. We also show examples from four countries, one from each of the four continents: Africa, America, Asia, and Europe.
For 36/39 countries, the native translated prompts yield worse performance compared to default English prompts. The exceptions are Indonesia, where Recall for the Indonesian native translated prompt is 1.0 higher than Recall for the English prompt, and Pakistan, where Recall for the Urdu native translated prompt is 0.7 higher than Recall for the English prompt.
The best-performing non-English language is often not the country’s language.
@@
We analyze the Recall scores for lower-income images from each of the 28 language prompts used in different countries. We observe that the best-performing language prompts for these countries are often not their own non-English major language. Specifically, for 24/39 countries, the best-performing non-English languages have better results than the default English prompts. However, surprisingly, in most cases, these languages are not spoken in these countries. The results are shown in Figure 3, where for 37/39 countries the language with the highest Recall (highlighted in yellow) is different from the non-English major language of the country (highlighted in cyan). The only exceptions are Indonesia and Pakistan, where the highest-performing languages and the non-English major languages are the same (highlighted in bold red).
Translated prompts decrease performance for all image income classes across all countries.
@@
We analyze the effect of the 40 non-English languages on all images from Dollar Street, regardless of which country the data was collected from. We include all 59 countries and results from all 40 translated prompts for this analysis. We aggregate Recall across all topics and countries for each set of non-English language prompts and group the scores according to image income classes. We measure the difference between Recall scores with default English prompts and the scores with translated prompts to determine the effect of each language on the four different image income classes.
@@
We show in Table 2 the average Recall and drops in performance for all the 40 translated prompts for each image income class. The most significant impact is on higher-income classes. Specifically, the rich and up-mid income classes experienced the highest drops in performance compared to other income classes. One potential reason is that images depicting rich and up-mid income classes are overrepresented in AI models and datasets, leading them to be seen as the “standard” representation for these topics. In a similar way, English is the most commonly used language for training AI models and is considered the “standard” language for textual data. Therefore, non-English, translated prompts may signal to the model that a representation different from the “standard” is being requested, leading to unpredictable, inferior results. We show the performance drops for each language in the Appendix Table 7.
@@
RQ2. Does adding country information improve performance for lower-income data?
@@
We calculate cosine similarity scores between the NLLB-CLIP-SigLIP image embeddings and the text embeddings of the 63 country suffix prompts. This results in 63 image-topic alignment scores for each image across all country suffixes. Similar to RQ1, as a baseline, we use the alignment scores between the images and the default English prompts.
Following the same procedure described in Section 4, we calculate the Recall scores for each topic using each of the 63 country suffix prompts. We analyze the effects of adding country suffixes to text prompts and present our findings below.
@@
Country-suffix prompts perform consistently better than default English prompts on lower-income images.
Given our focus for this analysis on low-income data, we filter out the 21 countries with no images from poor or low-mid income class households and focus on the remaining 42 countries.
@@
In Figure 4, we show the average Recall scores across all countries using the default English and country-suffix prompts. We also show Recall scores from four sample countries from different continents. From our results, we find that for most countries (38/42), adding the country suffix to the text prompts improves Recall performance for lower-income images compared to default English prompts. Countries where this does not hold are Bolivia, Brazil, Jordan, and the United States.
@@
The country-suffix prompt performance across different image income classes is influenced by the country’s economic status.
While country suffix prompts improve the performance of vision-language models on lower-income household data (poor), they simultaneously reduce performance on higher-income data.
@@
We measure the Recall of all data for each country suffix, group them into image income classes (based on household income), and place them into categories based on the World Bank income classification of that country suffix (e.g., Recall of data from poor households using country suffixes of poor countries is 31.2) For each image income class category, we calculate and show in Table 3 the average Recall of the country suffixes across the different country income levels and the average change in Recall compared to results obtained using the default English prompts. We show an extended table with results for all 63 country suffixes in the Appendix Table 8 and Table 9.
The differences in Recall between the country suffixes and default English prompts indicate which income group’s data benefits from an increase and vice versa. We find that country suffixes from countries that belong to the poor, low-mid, and up-mid World Bank Income categories lead to an increase in Recall for images with poor household income, as shown in green in Table 3. On the other hand, adding country suffixes to prompts leads to a decrease in Recall for images from the other household income groups (low-mid, up-mid, and rich).
@@
Interestingly, we observe that country suffixes tend to favor image retrieval from income groups that are the same or close to their World Bank income classification. The default English prompts favor the retrieval of up-mid income level data as the Recall for up-mid images is higher than the Recall for the other 3 income groups. We highlight in bold the average Recall that is highest among other income groups for each country suffix classification in Table 3, and show that poor country suffixes, low-mid country suffixes, and up-mid country suffixes produce the highest Recall scores for data from their corresponding income class. Rich country suffixes are the exception as they produce the highest Recall for data from the up-mid income level as opposed to the rich. We analyze the individual country suffix results and group lower-income images as images belonging to either the poor or lower-middle income group and higher-income images as images belonging to the upper-middle or rich income groups. We then find that 48/63 countries produce the highest Recall for lower-income data and are lower-income countries, or they have the highest Recall for higher-income data and are higher-income countries.
The best-performing country suffixes for lower-income data from a continent are from the same continent.
We analyze Recall results for lower-income data (from poor and low-mid households) from 42 countries when prompted using each of the 42 country suffixes and group these countries according to their respective continents. We categorize the data further into data from different World Bank Income classes within a continent. Then we display the average Recall and Recall difference (with respect to Recall using default English prompts) of each data group when prompted using country suffixes from different continents. For example, we show that lower-income data from poor African countries have a Recall of 36.6 and a performance increase of 15.7 when prompted using African country suffixes.
We find that the best-performing country suffixes for images from a continent are from the same continent, as shown by the diagonal of bold values in Table 4. We also see that lower-income data from African countries benefit most from the addition of country suffixes to prompts, while data from America and Asia do not experience Recall improvements.
Meanwhile, for higher-income data, there are no Recall improvements even when images and country suffixes belong to the same continent (see Appendix Table 10).
@@
RQ3. Does adding income information improve performance for lower-income groups?
We create three categories of income suffixes, poor, rich, and neutral, as described in Section 3.2. We repeat the image retrieval experiments from previous research questions to determine the Recall for images from each topic. We group and analyze the these results across countries and income groups.
@@
We find that the poor income suffix prompt obtains the best performance in 26/42 countries with lower-income images. For 12/42 countries, the default English prompts perform better than all the income suffixes. However, most countries experience Recall improvement from one of the income suffixes. We show the results in Figure 5. The average Recall score is aggregated across all 42 countries for the default English and income suffix prompts. The income suffix poor achieves the best recall. Recall scores for four countries is in Appendix Figure 7.
@@
We group the data into four income groups and further categorize them according to the World Bank income classification of the countries they were obtained. In Table 5, we show the Recall scores and performance increase with respect to the default English prompts for each data group.
We find that income suffixes benefit most data from poor households and some from lower-income families, while data from other income groups do not see an increase in Recall. In addition, there is no clear trend between the World Bank’s income classification of the country and the effect of income suffixes on Recall for the images. Further details about the income classification by World Bank for each country can be seen in Section A.2.

### v3

# Uplifting Lower-Income Data: Strategies for Socioeconomic Perspective Shifts in Large Multi-modal Models

@@ Abstract
Recent work has demonstrated that the unequal representation of cultures and socioeconomic groups in training data leads to biased Large Multi-modal(LMM) models. To improve LMM model performance on underrepresented data, we propose and evaluate several prompting strategies using non-English, geographic, and socioeconomic attributes. We show that these geographic and socioeconomic integrated prompts favor retrieving topic appearances commonly found in data from low-income households across different countries leading to improved LMM model performance on lower-income data. Our analyses identify and highlight contexts where these strategies yield the most improvements.
@@
Socioeconomic Perspective Shifts in Large Multi-modal Models
@@ 1 Introduction
A lack of diversity in popular AI datasets Shankar et al. (2017) leads to unequal model performance, further widening the technological gap between well-represented and underrepresented communities. While data from higher-income Western communities are readily available online, lower-income and non-Western data are often missing Rosling et al. (2019). As a result, cost-effective methods like web scraping fail to produce diverse datasets.
One approach to building large datasets leverages LMM models to filter uncurated data based on image-text association strength scores Fang et al. (2023). For instance, OpenAI’s ViT-B/32 Radford et al. (2021) was used to filter web-scraped images to create the LAION-5B datasetSchuhmann et al. (2022). However, foundation LMM models like CLIP perform unequally across cultures and socioeconomic groups, favoring higher-income and Western images Nwatu et al. (2023).
Datasets filtered by LMM models reflect the model’s biases Fang et al. (2023), often excluding underrepresented data and worsening the lack of diversity in AI models. Ignat et al. (2024) demonstrates this by showing that the LAION-5B dataset closely resembles data from Western countries, such as the United States and Canada while differing from non-Western countries’ data. This leads to LMM models with uneven performance on data drawn from different locations and income groups. Therefore, our paper seeks to answer the following question: How do we improve the performance of LMM models on lower-income and non-Western data?
We tackle performance inequality in LMM models Radford et al. (2021); Visheratin (2023) through prompting that transfers the cultural knowledge embedded in language Ventura et al. (2023); Buettner et al. (2024); Nguyen et al. (2024). Our goal is to improve the performance of LMM models on data from households with non-Western and lower socioeconomic status. Specifically, as shown in Fig. 1, we pose several research questions to evaluate the role of non-English languages, as well as prompts with geographic and socioeconomic attributes, to retrieve more diverse images.
Our contributions are summarized as follows. First, we show that a naive prompt translation-based approach fails to adequately address the performance gap of LMM models on lower-income data. Second, we establish that geographic and socioeconomic attribute integrated prompts improve LMM performance on lower-income data. We identify contexts where these prompts work best by conducting an in-depth analysis of LMM models’ understanding of these attributes and their effects on recall across data from different countries. Lastly, we share insights from our analysis demonstrating how these attributes drive a perspective shift that benefits the retrieval of lower-income data.
@@ 2 Related Work
Class imbalances in training data contribute significantly to bias in AI models Ferrara (2024); Shankar et al. (2017); He and Garcia (2009); Pouget et al. (2024), leading to unequal outcomes in areas like facial recognition Buolamwini and Gebru (2018), healthcare Obermeyer et al. (2019), and hiring Raghavan et al. (2020). Since creating balanced datasets is challenging and costly Ignat et al. (2024); Ramaswamy et al. (2023), researchers have explored bias mitigation techniques such as data augmentation, feature importance tuning, regularization, and adversarial training Yan et al. (2020); Zafar et al. (2017); Ignat et al. (2024); Maudslay et al. (2019); Sharma et al. (2020); Navarro et al. (2024); Zhang et al. (2018). Our work is most similar to research on post-processing methods Ferrara (2023); Hardt et al. (2016); Kamiran et al. (2012); Pleiss et al. (2017) that adjust model outcomes to meet diversity standards, aiming to benefit disadvantaged groups. Prior research has shown that LMM models perform poorly on data from lower socioeconomic groups, and our analysis investigates non-invasive post-processing methods to address this issue.
@@
Language plays a key role in transmitting cultural knowledge Callies (2024); Sharifian (2014); Karsdorp and Fonteyn (2019); Norton (1997), as AI models often absorb biases from the language in their training data Stanczak and Augenstein (2021); Rogers et al. (2021) and model outputs can be controlled by specifying a cultural shift in perspective Ventura et al. (2023) to improve diversity. However, research Arora et al. (2023); Cao et al. (2023); AlKhamissi et al. (2024); Liu et al. (2021) shows that large language models (LLMs) and LMM models capture more cultural information from English data (mainly Western) than from non-English data. This disparity stems from differences in the quantity and quality of non-English data, translation issues, and model design Arora et al. (2023); Hershcovich et al. (2022); Nasif et al. (1991).
@@
Recent studies have explored prompting techniques for large language models, including both hard Petroni et al. (2019); Zhou et al. (2023) and soft prompting Huang et al. (2023); Goswami et al. (2023), to improve model adaptation for tasks like instruction tuning, and value alignment. These methods are also applied in LMM models Lu et al. (2022); Yao et al. (2024); Zhou et al. (2022). While prior work Buettner et al. (2024) has incorporated geographic and physical attributes into prompts to enhance image retrieval diversity, this research extends the investigation to non-English language prompts and socioeconomic attributes to analyze how LMM models encode representations of various topics across regions and socioeconomic status.
@@ 3 Methodology
We propose prompting strategies that account for language, location, and socio-economic attributes and analyze how these prompts affect the performance of a multilingual LMM model on data across different socio-economic groups, primarily focusing on lower-income data.
@@ 3.1 Dollar Street Dataset
For fair comparison across bins, we group the images using the quartile binning method, which splits the data into an approximately equal number of images per bin as shown in Rojas et al. (2022). We group the images into four income classes (“poor”, “low-mid”, “up-mid”, and “rich” ) using quartiles as shown in Table 1. We further categorize the lowest two image income classes as lower-income images and the highest two income groups as higher-income images.
@@
Country Economic Classes.
We group all 63 countries from Dollar Street into country economic classes based on their World Bank income classification.111https://datahelpdesk.worldbank.org/ All the countries and their economic classes are shown in Section A.1. We further categorize the lowest two country economic classes as lower-income countries and the highest two economic groups as higher-income countries.
@@
There are unique topics associated with the images in the dataset which reflect everyday household objects and human actions (e.g., “toilet paper”, “get water”), some of which are subjective (e.g., “next big thing I plan to buy”, “favorite sports clubs”, “most loved item”). We remove nineteen subjective topics from the dataset following De Vries et al. (2019) and Nwatu et al. (2023).
@@ 3.2 Prompt Design
We describe below the prompting strategies we use for our experiments and show examples in Figure 1.
@@
Specifically, we pair each country with their non-English major language (e.g., Portuguese for Brazil, French for Cameroon) following the country and language information provided by official sources.222www.cia.gov/the-world-factbook/field/languages/, www.ncsc.org/__data/assets/pdf_file/0024/17862/languagesbycountries.pdf, www.dss.gov.au/sites/default/files/files/foi_disclosure_log/12-12-13/language-list.pdf
We identify 59/63 countries in Dollar Street where one or more major non-English languages are spoken. We also select languages covered by state-of-the-art machine translation and multilingual LMM models. There are 40 such non-English major languages, and they are listed in Section A.1.
Finally, we translate the default English prompts to these 40 languages using the NLLB-200-distilled-600M Costa-jussà et al. (2022), an open-source state-of-the-art neural machine translation model. Translation metrics for NLLB-200-distilled-600M are shown in Appendix Table 13 and available on HuggingFace. If an image prompt is translated into the non-English major language of the image’s country of origin, it is referred to as a native translated prompt.
@@
3.3 State-of-the-art LMM Model
For our evaluation, we chose NLLB-CLIP-SigLIP Visheratin (2023), a state-of-the-art multilingual LMM model, due to its broad reach across many low-resource languages and superior performance among other models.333https://huggingface.co/visheratin/nllb-clip-large-siglip The model consists of an image encoder from the SigLIP model Beyer et al. (2022); Zhai et al. (2023) and a text encoder from the NLLB model Costa-jussà et al. (2022). The model supports the 201 languages of the Flores-200 Costa-jussà et al. (2022) and has recorded groundbreaking results on the Crossmodal-3600 dataset Thapliyal et al. (2022), especially on low-resource languages.
@@ 4 Research Questions
We perform several analyses to answer three research questions that uncover and mitigate limitations in the performance of LMM models across different countries and socioeconomic groups.
@@
4.1 RQ1. Do translated prompts improve retrieval performance for lower-income images?
@@
We calculate the cosine similarities between image and translated prompt text embeddings for each image-topic pair across English and 40 non-English languages, generating 41 alignment scores per image. The alignment scores with default English prompts serve as our baseline.
@@
We compute Recall scores by selecting the top N images with the highest alignment scores for each topic, where N represents the number of ground truth images. We then group and analyze the Recall scores across different countries and image income classes and present our findings below.
Native translated prompts perform
consistently worse than English prompts on
lower-income images from their respective countries.
@@
We focus our analysis on images from the two lowest image income groups, i.e., poor and low-middle as grouped in 3.1. After excluding 20 countries without data for these income groups (e.g., Russia, Turkey), we retain 39/59 countries and 28/40 non-English languages for the study. Each country is paired with its native non-English language, and we compare Recall scores for the native translated prompts to those for the default English prompts. The average Recall across all countries and scores from four countries are displayed in Figure 2.
For 35 out of 39 countries, the native translated prompts underperform compared to the default English prompts. The exceptions include Burkina Faso, Nigeria, Pakistan and Tanzania, where native translated prompts in French (diff. of 1.0), Hausa (diff. of 0.2), Urdu (diff. of 0.7) and Swahili (diff. of 1.5), respectively, outperform English prompts. Overall, native translated prompts generally fail to retrieve diverse images, as depicted by the example using French prompt in Fig. 1.
The best-performing non-English language
often differs from the country’s native
language.
@@
We analyze the Recall scores for lower-income images across 28 language prompts used in different countries and find that the best-performing language prompts often differ from the countries’ non-English major languages. Specifically, in 24 out of 39 countries, non-English language prompts outperform the default English prompts, yet these top-performing languages are not typically spoken in the respective countries. As illustrated in Figure 3, for 37/39 countries, the language with the highest Recall score (highlighted in yellow) differs from the country’s primary non-English language (highlighted in cyan), with exceptions in Indonesia and Pakistan, where they coincide (highlighted in bold red).
Translated prompts decrease performance for
all image income classes across all
countries.
@@
We analyze the impact of 40 non-English language prompts on all images from Dollar Street, covering 59 countries and we group Recall scores by image income classes. By comparing Recall scores between default English prompts and native translated prompts, we assess the effect of each non-English language on four income classes and show the difference in scores in Appendix Table 10.
@@
We show in Table 2 the average Recall and drops in performance across all 40 translated prompts for each image income class. The results indicate that higher-income classes, specifically the rich and up-mid groups, experience the largest drops in performance with translated prompts. This may be due to the overrepresentation of images from these income groups in AI models and datasets, positioning them as the "standard" representation. Similarly, English, the dominant training language, is seen as the “standard” for textual data, so non-English prompts may signal a deviation from this standard, resulting in poorer model performance.
@@
4.2 RQ2. Does adding country information improve retrieval performance for lower-income images?
@@
We compute cosine similarity scores between NLLB-CLIP-SigLIP image embeddings and the text embeddings of 63 country suffix prompts., yielding 63 image-topic alignment scores per image. Using the alignment scores from the default English prompts as a baseline, we follow the procedure outlined in Section 4.1 to calculate Recall scores for each topic with the country suffix prompts. We then analyze the impact of adding country suffixes to text prompts and present the results in the following sections.
@@
Country-suffix prompts perform consistently better than default English prompts on lower-
income images.
Focusing on low-income data, we filter out 21 countries without images from poor or low-mid income households, leaving 42 countries for analysis. In Figure 4, we present the average Recall scores across all countries using both default English and country-suffix prompts, along with results from four sample countries from different continents.
@@
Our findings indicate that in 38/42 countries, adding a country-suffix to text prompts improves Recall performance for lower-income images compared to default English prompts. Exceptions include Bolivia, Brazil, Jordan, and the United States. Country-suffix prompts are thus more effective in retrieving diverse images, as demonstrated by the Cameroon example in Fig. 1.
@@
A country’s economic status influences the performance of its country-suffix prompt across different image income classes.
@@
Country suffix prompts improve LMM model performance for lower-income images (poor) but reduce performance for higher-income images. Using World Bank income classifications, we calculate Recall scores across four country suffix groups (poor, low-mid, up-mid, and rich) and four image income classes (based on household income). For each image income class, we aggregate Recall scores and compare them with those from default English prompts, as shown in Table 3, with detailed results in Appendix Table 11 and Table 12. For example, Table 3 shows that Recall of images from poor households using country suffixes of poor countries is 31.2, a 9.7 increase from default English prompt performance on that group.
The analysis reveals that country suffixes from poor, low-mid, and up-mid income categories improve Recall for images from poor households, while reducing Recall for higher-income groups (low-mid, up-mid, rich).
@@
Interestingly, country suffixes tend to favor image retrieval from income groups that match or are close to their own economic classification. When the four income classes are re-categorized into two classes (lower-income: poor, low-mid and higher-income: up-mid, rich), we find that in 48/63 cases, the image income category with the highest Recall corresponds to the country suffix’s economic class, demonstrating the alignment between income levels of country-suffixes and retrieval performance.
The best-performing country suffixes for lower-income images from a continent are from the same continent.
We calculate Recall results for lower-income images from 42 countries using the 42 country suffix prompts, yielding a total of 1,764 Recall scores. Using country-suffixes, we group these scores by continent and further categorize them based on the World Bank Income classes of the respective country-suffixes. We present the average Recall and differences compared to default English prompts for each group in Table 4.
@@
Our findings emphasize the significance of regional specificity in data collection, as the best-performing suffixes align with their respective continents (shown by the diagonal of bold values in Table 4). The results indicate that lower-income images from African nations benefit significantly from including country suffixes. In contrast, data from America and Asia show no Recall improvements, underscoring the necessity for tailored data collection strategies across regions. Notably, lower-income images from African countries exhibit a Recall score of 36.6, reflecting the highest performance increase of 15.7 when using African suffixes. The positive impact of country suffix prompt additions is particularly pronounced for Africa, as the prompts enhance performance on underrepresented data by shifting model inference from its learned standard. This effect is crucial given the current datasets often lack representation from African countries and poor households. Additionally, the similarities among African countries contribute to this improved performance.
Meanwhile, we find no Recall enhancements for higher-income data, regardless of the alignment between images and country suffixes (see Table 5).
| Image by Continent | Country by Income | Country Suffix | |||
| Africa | America | Asia | Europe | ||
| Africa | Poor | 42.3 (-1.4) | 40.3 (-3.4) | 36.0 (-7.7) | 40.3 (-3.4) |
| Low-mid | 41.8 (-8.8) | 39.7 (-10.9) | 37.0 (-13.6) | 40.7 (-9.9) | |
| Up-mid | 33.3 (-5.9) | 32.4 (-6.8) | 25.7 (-13.5) | 33.4 (-5.8) | |
| Average | 39.1 (-5.4) | 37.5 (-7.0) | 32.9 (-11.6) | 38.1 (-6.4) | |
| America | Up-mid | 24.3 (-20.8) | 37.6 (-7.5) | 26.9 (-18.3) | 37.2 (-8.0) |
| Rich | 28.7 (-33.4) | 45.7 (-16.4) | 32.5 (-29.6) | 44.7 (-17.4) | |
| Average | 26.5 (-27.1) | 41.7 (-12.0) | 29.7 (-24.0) | 41.0 (-12.7) | |
| Asia | Low-mid | 33.8 (-17.2) | 38.4 (-12.6) | 40.4 (-10.5) | 43.4 (-7.6) |
| Up-mid | 30.9 (-21.0) | 40.5 (-11.4) | 41.2 (-10.7) | 46.5 (-5.4) | |
| Rich | 28.2 (-19.9) | 36.1 (-12.0) | 36.5 (-11.6) | 40.1 (-8.0) | |
| Average | 31.0 (-19.4) | 38.3 (-12.0) | 39.4 (-10.9) | 43.3 (-7.0) | |
| Europe | Low-mid | 22.3 (-23.9) | 33.3 (-13.1) | 26.6 (-19.6) | 42.2 (-0.4) |
| Up-mid | 18.9 (-25.0) | 29.2 (-14.7) | 24.5 (-19.4) | 40.7 (-3.2) | |
| Rich | 19.0 (-20.8) | 27.8 (-12.0) | 21.5 (-18.3) | 37.6 (-2.2) | |
| Average | 20.1 (-23.2) | 30.0 (-13.3) | 24.2 (-19.1) | 40.2 (-3.1) | |
4.3 RQ3. Does adding income information improve retrieval performance for lower-income images?
We create three categories of income suffixes, poor, rich, and neutral, as described in Section 3.2. We repeat the image retrieval experiments from previous research questions to determine the Recall for images from each topic. We group and analyze these results across countries and income groups.
@@
Our analysis reveals that the poor income suffix prompt achieves the highest performance in 26/42 countries with lower-income images. In 12/42, default English prompts outperform all income suffixes. Nevertheless, most (30/42) countries show Recall improvements when using one of the income suffixes.
We illustrate in Figure 5 the aggregate the average Recall scores for all 42 countries across default English and the income suffix prompts. Notably, the poor income suffix demonstrates the best Recall, effectively retrieving a diverse array of images, as shown by the example in Fig. 1. Recall scores for four sample countries are in Appendix Figure 7.
@@
We group the data into four income groups (by household income) and further categorize them according to the World Bank income classification of their country of origin. In Table 6, we show the Recall scores and performance improvements relative to the default English prompts for each data group.
We find that income suffixes predominantly benefit data from poor households and some from low-mid income households, while data from other income groups do not show Recall increases.